OpenStudy (anonymous):

A line has a slope of -1.5 and a y-intercept of (0,3). What is the value of y when x = 6 ? @iambatman @Hero @zephyr141 @zepdrix @linn99123 @mathmath333

OpenStudy (zale101):

y = mx + b where m is the slope -1.5 and b is the y-intercept (0,3) plug those in and create your linear equation. Afterwards plug in x=6 to find y.

OpenStudy (anonymous):

-1.5 * 6 = 9 + 3 =12 y = 12?

OpenStudy (anonymous):

@Zale101 ?

OpenStudy (anonymous):

Wait, it would be -9. So -9 + 3 = -6

OpenStudy (zale101):

y = mx + b y = (-1.5)x + (3) <-- your linear equation y=(-1.5)(6)+3 <--- plug in x=6 and solve for y

OpenStudy (anonymous):

y = -6 ?

OpenStudy (zale101):

yes

OpenStudy (anonymous):

thanks!

OpenStudy (zale101):

You're welcome :)